Solve
9x^2 - 12x - 5

Thank You

Answer by Nate(3500)

I take it that the expression you showed is equal to zero:
9x^2 - 12x - 5 = 0 find factors of 9 * -5 = -45 that sum to -12 ~~~> -15,3
9x^2 - 15x + 3x - 5 = 0
3x(3x - 5) + 1(3x - 5) = 0
(3x + 1)(3x - 5) = 0
3x + 1 = 0
3x = -1
x = -1/3
and
3x - 5 = 0
3x = 5
x = 5/3